# Fear over Friends: Examining the Perceived Influence of Others on Vaccination Decisions

### v1

@@ 3 Results
In Figure 3, we show the distributions of political leaning, rural-urban identity, and education. These demographics and others are included in our ordinal logistic regression analysis found in Table 1. Additionally, in Figure 1, we show a map of the U.S. counties in which survey participants resided in at the time of the survey.
Across both the vaccinated and unvaccinated groups, participants described themselves as 40.7% male, 57.01% female, and 2.02% non-binary/third gender, and 86.1% white, 6.8% black, 3.5% Asian, 2.8% other, and 0.8% Native American. There were no significant differences found between these demographics across the two groups.
@@
There were significant differences across other demographic traits, such as political leaning, rural-urban identity, rural residency, income, and education. Specifically, vaccinated participants were significantly more left leaning, identified more as city/suburban people, lived in more urban areas, had higher income, and more education.
@@
These demographic differences align closely with results from previous work. For example, reports from The COVID States project [^2] https://www.covidstates.org/ found that people who lived in city/suburban areas, had more education, and had higher income, had significantly higher vaccination rates and greater support for vaccine requirements lazer2021covid ; baum2021covid . Other studies have shown that rural residents are “significantly less likely to participate in COVID-19-related preventive health behaviors callaghan2021rural ” and that rural residents have higher rates of vaccine resistance lazer2021covid . Further, as stated by Green et al. (2022), “partisanship remain[s] the most stable and sizable gap” in COVID-19 behaviors and attitudes green2022using . Prior work has shown that vaccine hesitancy is significantly higher for those who identify as Republications than those who identify as Democrats reiter2020acceptability ; perlis2020covid ; khubchandani2021covid ; milligan2021covid ; clinton2021partisan . The demographic results from our survey provide further robustness to these prior findings.
Next, we examine our core survey questions capturing the perceived influence of others on vaccine decisions and the perceived agreement of others with vaccine decisions. In Table 1 we show the results of our ordinal logistic regression analysis over the perceived influence and agreement across theoretically strong and weak social ties.
At a high level, no matter the social group, vaccinated participants perceived significantly higher influence from others and significantly higher agreement with others than participants who were not vaccinated. For example, from Table 2, given that the coefficient estimates from our regression analysis are given in units of ordered logits or ordered log odds, we can say that for one unit increase in the independent variable (going from no to yes for the question “Have you received one or both shots for COVID-19?”), we expect a 1.0974 increase in the expected value of perceived influence of close friends (ordered as a 5-point scale from None at all to A great deal), meaning that vaccinated participants perceived significantly more influence from their close friends than participants who are unvaccinated. Similarly, we can say that we expect a 2.2421 increase in the value of perceived agreement of close friends with the participants vaccination decisions (ordered as a 3-point scale from disagree to agree), indicating vaccinated participants perceived that their close fiends agreed with their decision to get vaccinated significantly more than participants who were not vaccinated (i.e. close friends disagree with their decision to not get vaccinated). The only exception to this finding is the perceived influence of one’s ‘Church or Social club’ - which showed no significant differences between the vaccinated and not vaccinated groups. Further, the perceived influence and agreement weakens slightly (although not consistently) as we move from theoretically stronger relationships to weaker relationships (e.x. family to neighbors), yet statistically significant differences between the vaccination groups still remain across the social circles.
These findings align with the findings of previous work. Tunçgenç et al. (2021) showed that the perceived approval and adherence of others to social distancing during COVID-19 predicted participant’s adherence to social distancing, particularly when others within one’s close social circle approved and adhered to pandemic guidelines tunccgencc2021social . However, this study did not demonstrate differences in perceived social influence between groups, rather the authors demonstrated that there is a relationship between individual adherence and perceptions of social circle adherence. We add nuance to this finding, by showing perceptions of influence and approval may not be symmetrically across groups (those who follow pandemic guidelines and those who do not).
@@
Next, we examine how the perceived danger of COVID-19 to varying social groups changes across vaccinated and unvaccinated participants. Specifically, we ask: “How likely do you think members of each social group are to get severe COVID-19 or die from COVID-19?” - capturing an indirect social influence rather than a direct influence. It is reasonable to assume that if one believes COVID-19 is dangerous to their family and close friends, they may be more likely to get vaccinated to protect them.
We find similarly divided results between the two groups. As shown in Table 2, we find that across all social groups vaccinated participants perceived COVID-19 as more dangerous to others than unvaccinated participants. While there are some variations across social groups (i.e. perceiving family as in more danger than close friends, etc.), those variations are not significantly different.
@@
While to the best of our knowledge, perceived danger of COVID-19 to specific social groups has not been studied, our results align with more general findings in the literature. Many factors have been shown to predict COVID-19 risk perceptions, such as personal experience with the virus and hearing about the virus from friends and family dryhurst2020risk . Hearing about the virus through friends and family was shown to be the strongest predictor of perceived risk, where risk includes both risk of self and risk of others contracting the virus dryhurst2020risk . Other non-social factors have also been shown to predict personal risk to COVID-19, including psychological factors and institutional trust dryhurst2020risk ; yildirim2022factor .
(a) Family (b) Close Friends
[Uncaptioned image] [Uncaptioned image]
(c) Co-workers (d) Neighbors
[Uncaptioned image] [Uncaptioned image]
[…]
figureDistributions of perceived influence on vaccination decision. Due to space restrictions, we leave out the plot for “Church or Social Club”. Note the changes in the x-axis across the social groups.
Demographics
Ordinal Question Coef 95% CI Interpretation
Political Leaning -1.94*** -2.19 to -1.69 More liberal
Rural-Urban Identity -0.64*** -0.87 to -0.41 More identify as city people
Rural-Urban Residency -0.44*** -0.68 to -0.21 Live in more urban areas
Income 0.63*** 0.42 to 0.85 Higher income
Education 0.95*** 0.72 to 1.18 More educated
Perceived Influence
Ordinal Question Coef 95% CI Interpretation
Family 1.07*** 0.84 to 1.31 More influenced
Close friends 1.10*** 0.84 to 1.35 More influenced
Co-workers 1.07*** 0.73 to 1.40 More influenced
Church or Social club 0.38 -0.06 to 0.82 No difference
Neighbors 1.22*** 0.86to 1.59 More influenced
Perceived Agreement
Ordinal Question Coef 95% CI Interpretation
Family 1.93*** 1.64 to 2.22 More agreement
Close friends 2.24*** 1.95 to 2.53 More agreement
Co-workers 1.91*** 1.62 to 2.21 More agreement
Church or Social club 1.05*** 0.77 to 1.33 More agreement
Neighbors 1.77*** 1.46 to 2.07 More agreement
Perceived Danger of COVID-19
Ordinal Question Coef 95% CI Interpretation
Family 0.68*** 0.46 to 0.91 More dangerous
Close friends 0.49*** 0.26 to 0.71 More dangerous
Co-workers 0.66*** 0.43 to 0.89 More dangerous
Church or Social club 0.83*** 0.60 to 1.07 More dangerous
Neighbors 0.84*** 0.61 to 1.07 More dangerous
Table 1: Results from 20 ordinal regression models, where being vaccinated is the independent variable. If the 95% CI does not cross 0, the parameter estimate is statistically significant, also shown by the p-value. Note, that *** indicates $p<0.001$ and ** indicates $p<0.01$. The coefficient estimates provided are given in units of ordered logits, or ordered log odds. So, for example, for the independent variable - ‘Have you received one or both shots for COVID-19?’ and the dependent variable ‘What is your political leaning?’- we would say that for a one unit increase in the IV (i.e., going from ‘No’ to ’Yes’), we expect a $-1.94$ decrease in the expected value of political leaning (ordered as Very Liberal to Very Conservative) on the log odds scale, meaning vaccinated participants were more politically liberal. The interpretation column indicates a trait of vaccinated participants.
@@
We can further contextualize our findings by approximating the size of and trust in participants social groups. While theoretically, our survey questions measure differences across strong and weak social ties, the strength of the ties between participants and these groups likely changes across individuals. Hence, here we utilize the answers to the questions asked about voluntary conversations participants had within the past week and the county in which participants lived during the pandemic, as described in the methods section above.
@@
To better measure differences between the two groups, we perform a Mann-Whitney test on each pair of distributions, finding significant differences between approximate social circle size (pvalue=1.6301e-07) and the number of individuals trusted (pvalue=1.8238e-05), but no significant difference was found between the proportion of social circle members trusted (pvalue=0.9807). Hence, vaccinated participants interacted with slightly more people than unvaccinated, but trust in those people is no different between the groups.
As another proxy for the number of people participants encounter, we can also look at the areas of the U.S. in which participants live. Overall, 51.62% of participants, resided in metro areas with 1 million population or more (RUCC 1). More broadly, 86.36% of participants lived in metro counties (RUCCs 1 through 3) and only 13.64% lived in nonmetro counties (RUCCs 4 through 9). Of the vaccinated participants, 57.25% of participants, resided in metro areas of 1 million population or more, 89.02% lived in metro counties, and 10.98% lived in nonmetro counties. Of the unvaccinated participants, 45.63% of participants, resided in metro areas of 1 million population or more, 83.54% lived in metro counties, and 16.46% lived in nonmetro counties.
As shown in Table 1, when examining rural-urban residency using an ordered logistic regression model, we find that the vaccinated participants live in more urban areas than unvaccinated participants. Precisely, for one unit increase in the independent variable (going from no to yes for the question “Have you received one or both shots for COVID-19?”), we expect a -0.4429 decrease in the expected value of RUCC (ordered as a 9-point scale from 1 to 9), which is statistically significant.
Hence, both the approximations of the number of people participants interacted with prior to the survey (social circle size and residency) suggest that the vaccinated population interacted with more people than the unvaccinated population did. Theoretically, it makes sense that interacting with less people can impact one’s perceptions of COVID-19, as interacting with less people reduces the chance that one comes in close contact with someone who had been severely ill from COVID-19. In general, this idea is supported in the literature. For example, when studying vaccine messaging strategies, it was found that “messages evoking harm reduction and people you know were more effective in counties where the virus is spreading more quickly. green2021covid
There are likely other confounding factors related to location that may have a significant impact on one’s pandemic-related decisions. For example, prior work has demonstrated that local news coverage during 2020 and 2021 was best explained by national trends rather than by local conditions, and the themes of that coverage varied across counties in the U.S. joseph2022local . Further, we know that there were urban-rural differences in COVID-19 behaviors, and that those behaviors changed depending on the news produced locally. Specifically, Kim et al. (2020) showed that “rural residents [were] more likely to engage in social distancing behavior than otherwise similar rural residents if their local news [was] produced in a city that is more impacted by COVID-19” kim2020effect . Other work has shown wide variations in vaccine support across areas in the United States perlis2020covid .

### v2

@@ 3.1 Demographic Differences
Before examining our research questions, we describe the demographic differences across the vaccinated and unvaccinated groups. In Figure 1, we show the correlation among our key demographic variables and distributions of political leaning, rural-urban identity, and education across vaccinated and unvaccinated groups. Additionally, in Figure 1a, we show a map of the U.S. counties in which survey participants resided in at the time of the survey. In Table 1, we show the results of an ordinal logistic regression mdoel, where each demographic question is an independent variable and whether a participant was vaccinated or not is the dependent variable.
coef std err z P>|>|> |z|||| [0.025 0.975]
Education 0.3130 0.058 5.429 0.000*** 0.200 0.426
Income 0.1218 0.025 4.963 0.000*** 0.074 0.170
Political Leaning 1.0445 0.080 13.126 0.000*** 0.889 1.200
Rural-Urban Identity 0.1310 0.085 1.534 0.125 -0.036 0.298
Rural-Urban Residency 0.0166 0.048 0.348 0.728 -0.077 0.110
0.0/1.0 5.7225 0.474 12.074 0.000*** 4.794 6.651
Table 1: Ordinal logistic regression results for key demographic variables. The DV is a binary variable of which group a participant is in, where 1 is for vaccinated and 0 is for unvaccinated. Significance codes are shown for p: *** $p<0.001$, ** $p<0.01$, * $p<0.05$.
@@
Across both the vaccinated and unvaccinated groups, participants described themselves as 40.7% male, 57.0% female, and 2.0% non-binary/third gender, and 86.1% white, 6.8% black, 3.5% Asian, 2.8% other, and 0.8% Native American. There were no significant differences found between gender or race across the two groups. There were significant differences across other demographic traits, such as political leaning, income, and education. Specifically, vaccinated participants were significantly more left leaning, had higher income, and were more educated. When examining demographic distributions independently, we found some evidence that there were differences between rural-urban identity and rural residency; namely, vaccinated participants identified more as city/suburban people and lived in more urban areas (shown in the Supplemental materials). However, when controlling for the differences in political leaning, income, and education, these differences become non-significant.
@@
These demographic differences align with results from previous work. For example, reports from The COVID States project [^1] https://www.covidstates.org/ found that people who lived in city/suburban areas, had more education, and had higher income, had significantly higher vaccination rates and greater support for vaccine requirements lazer2021covid ; baum2021covid . Other studies have shown that rural residents are “significantly less likely to participate in COVID-19-related preventive health behaviors callaghan2021rural ” and that rural residents have higher rates of vaccine resistance lazer2021covid . Further, as stated by Green et al. (2022), “partisanship remain[s] the most stable and sizable gap” in COVID-19 behaviors and attitudes green2022using . Prior work has shown that vaccine hesitancy is significantly higher for those who identify as Republications than those who identify as Democrats reiter2020acceptability ; perlis2020covid ; khubchandani2021covid ; milligan2021covid ; clinton2021partisan . The demographic results from our survey provide further robustness to these prior findings as well as validate our survey data.
3.2 Perceived Social Influence (RQ1)
coef std err z P>|>|> |z|||| [0.025 0.975]
Education 0.3430 0.062 5.559 0.000*** 0.222 0.464
Income 0.1075 0.026 4.112 0.000*** 0.056 0.159
Political Leaning 1.0386 0.084 12.321 0.000*** 0.873 1.204
Rural-Urban Identity 0.0966 0.089 1.089 0.276 -0.077 0.271
Rural-Urban Residency 0.0073 0.050 0.147 0.883 -0.091 0.106
Family Influence 0.3097 0.077 4.014 0.000*** 0.158 0.461
Close Friends Influence 0.2184 0.104 2.095 0.036* 0.014 0.423
Coworkers Influence 0.2119 0.148 1.429 0.153 -0.079 0.502
Church/Club Influence -0.3445 0.212 -1.626 0.104 -0.760 0.071
Neighbors Influence 0.4425 0.195 2.268 0.023* 0.060 0.825
0.0/1.0 7.0613 0.576 12.253 0.000*** 5.932 8.191
Table 2: Ordinal logistic regression results for perceived social influence controlling for key demographic variables. The DV is a binary variable of which group a participant is in, where 1 is for vaccinated and 0 is for unvaccinated. Significance codes are shown for p: *** $p<0.001$, ** $p<0.01$, * $p<0.05$.
Figure 2: Distributions of perceived influence across social groups of vaccinated (blue) and unvaccinated (yellow) participants. The means perceived influence across each social group, where ‘None at all’ is 1 and ‘A great deal’ is 5, where: (a) 2.57 for vaccinated, 1.74 for unvaccinated, (b) 2.11 for vaccinated, 1.47 for unvaccinated, (c), 1.47 for vaccinated, 1.17 for unvaccinated, (d) 1.15 for vaccinated, 1.12 for unvaccinated, (e) 1.40 for vaccinated, 1.12 for unvaccinated.
First, we examine differences in perceived social influence of theoretically strong and weak social ties across vaccinated and unvaccinated participants. In Table 2, we again show the results of fitting an ordinal logistic regression model, where the dependent variable is if the participant is vaccinated or not. However, now we add our influence variables to the model (i.e., “How much influence did each social group have on your decision to (not) get vaccinated?”). This analysis allows us to show differences in perceived influence relative to other social groups and control for demographic differences. In Figure 2, we show the distributions of perceived influence across each social circle in terms of $N$ per Likert scale answer per vaccination group.
What is clear from this analysis is that vaccinated participants perceived significantly higher influence from their family than unvaccinated participants ($p<0.001$). Vaccinated participants also perceived significantly more social influence from close friends and neighbors (each $p<0.05$), but these differences were less than the perceived influence of family. On average, vaccinated participants fell between ‘a little’ and ‘a moderate amount’ for family influence ($\mu=2.57$, $\sigma=1.48$), between ‘a little’ and ‘a moderate amount’ for close friends influence ($\mu=2.11$, $\sigma=1.30$), and between ‘None at all’ and ‘A little’ for neighbor influence ($\mu=1.40$, $\sigma=0.81$). On the other hand, unvaccinated participants fell between ‘None at all’ and ‘A little’ for family influence ($\mu=1.74$, $\sigma=1.02$), close friends influence ($\mu=1.46$, $\sigma=0.88$), and neighbor influence ($\mu=1.12$, $\sigma=0.44$). Despite some significant differences across these groups, the majority of participants from both the vaccinated and unvaccinated groups did not perceive any social influence across all social ties (as shown in Figure 2). There were no significant differences in the perceived influence of coworkers or church/club members. When examining the distributions in Figure 2, we see generally the same trend across both groups. That is, perceived influence by both groups generally weakens as we move from theoretically stronger relationships to weaker relationships (e.x. family to church/club).
3.3 Perceived Social Agreement (RQ2)
coef std err z P>|>|> |z|||| [0.025 0.975]
Education 0.2442 0.070 3.494 0.000*** 0.107 0.381
Income 0.1174 0.030 3.974 0.000*** 0.060 0.175
Political Leaning 1.0520 0.094 11.216 0.000*** 0.868 1.236
Rural-Urban Identity 0.0931 0.098 0.952 0.341 -0.099 0.285
Rural-Urban Residency 0.0372 0.054 0.694 0.488 -0.068 0.142
Family Agree 0.7953 0.143 5.555 0.000*** 0.515 1.076
Close Friends Agree 1.0437 0.176 5.924 0.000*** 0.698 1.389
Coworkers Agree 0.5954 0.202 2.951 0.003** 0.200 0.991
Church/Club Agree -0.6762 0.230 -2.937 0.003** -1.127 -0.225
Neighbors Agree 1.2664 0.210 6.042 0.000*** 0.856 1.677
0.0/1.0 12.5360 0.838 14.953 0.000*** 10.893 14.179
Table 3: Ordinal logistic regression results for perceived social agreement controlling for key demographic variables. The DV is a binary variable of which group a participant is in, where 1 is for vaccinated and 0 is for unvaccinated. Significance codes are shown for p: *** $p<0.001$, ** $p<0.01$, * $p<0.05$.
Figure 3: Distributions of perceived agreement across social groups of vaccinated (blue) and unvaccinated (yellow) participants. The means perceived agreement across each social group, where ‘Disagree’ is 1 and ‘Agree’ is 3, where: (a) 2.78 for vaccinated, 2.19 for unvaccinated, (b) 2.80 for vaccinated, 2.19 for unvaccinated, (c), 2.48 for vaccinated, 1.98 for unvaccinated, (d) 2.21 for vaccinated, 1.95 for unvaccinated, (e) 2.34 for vaccinated, 1.88 for unvaccinated.
Next, we examine the differences between vaccinated and unvaccinated participants in perceived agreement from varying social circles. In Table 3, we show the results of fitting an ordinal logistic regression model, where the dependent variable is if the participant is vaccinated or not and a mix of demographic controls and perceived agreement are independent variables (i.e., “How much do you think each social group agrees with your decision to (not) get vaccinated?”). In Figure 3, we show the distributions of perceived agreement across varying social circles.
At a high level, no matter the social group, vaccinated participants perceived significantly higher agreement with others than participants who were not vaccinated. Perceived agreement from family ($p<0.001$), close friends ($p<0.001$), coworkers ($p<0.01$), and neighbors ($p<0.001$) all were perceived as higher for vaccinated participants than unvaccinated participants. The only exception to this finding is the perceived influence of one’s ‘Church or Social club’. On average, vaccinated participants reported perceived agreement between ’Do not care’ and ’Agree’ for family ($\mu=2.78$, $\sigma=0.53$), close friends ($\mu=2.80$, $\sigma=0.44$), coworkers ($\mu=2.48$, $\sigma=0.57$), neighbors ($\mu=2.34$, $\sigma=0.57$), and church/club ($\mu=2.20$, $\sigma=0.56$). On average, unvaccinated participants reported perceived agreement between ‘Do not care’ and ‘Agree’ for family ($\mu=2.19$, $\sigma=0.76$) and close friends ($\mu=2.20$, $\sigma=0.63$), although closer to ‘Do not care’ and to ’Agree’. Further, unvaccinated participants perceived agreement between ‘Disagree’ and ‘Do not care’ for coworkers ($\mu=1.99$, $\sigma=0.50$), neighbors ($\mu=1.88$, $\sigma=0.53$), and church/club ($\mu=1.94$, $\sigma=0.52$). The same general trend of weakening as we move from theoretically stronger relationships to weaker relationships that we saw for perceived influence also occurs for perceived agreement, particularly for the vaccinated group.
3.4 Perceived Social Danger (RQ3)
coef std err z P>|>|> |z|||| [0.025 0.975]
Education 0.3236 0.059 5.468 0.000*** 0.208 0.440
Income 0.1391 0.025 5.487 0.000*** 0.089 0.189
Political Leaning 0.9763 0.082 11.953 0.000*** 0.816 1.136
Rural-Urban Identity 0.1153 0.086 1.334 0.182 -0.054 0.285
Rural-Urban Residency 0.0082 0.048 0.169 0.866 -0.087 0.103
Family Danger 0.1300 0.096 1.356 0.175 -0.058 0.318
Close Friends Danger 0.0447 0.119 0.377 0.706 -0.188 0.277
Coworkers Danger 0.0152 0.129 0.118 0.906 -0.237 0.268
Church/Club Danger 0.0923 0.125 0.740 0.459 -0.152 0.337
Neighbors Danger 0.1904 0.126 1.506 0.132 -0.057 0.438
0.0/1.0 6.7926 0.543 12.510 0.000*** 5.728 7.857
Table 4: Ordinal logistic regression results for perceived social danger controlling for key demographic variables. The DV is a binary variable of which group a participant is in, where 1 is for vaccinated and 0 is for unvaccinated. Significance codes are shown for p: *** $p<0.001$, ** $p<0.01$, * $p<0.05$.
Figure 4: Distributions of perceived danger across social groups of vaccinated (blue) and unvaccinated (yellow) participants. The means perceived danger across each social group, where ‘Extremely unlikely’ is 1 and ‘Extremely likely’ is 5, where: (a) 2.70 for vaccinated, 2.28 for unvaccinated, (b) 2.28 for vaccinated, 1.97 for unvaccinated, (c), 2.55 for vaccinated, 2.17 for unvaccinated, (d) 2.82 for vaccinated, 2.36 for unvaccinated, (e) 2.90 for vaccinated, 2.42 for unvaccinated.
@@ 3.4 Perceived Social Danger (RQ3)
Lastly, we examine how the perceived danger of COVID-19 to varying social groups changes across vaccinated and unvaccinated participants. Specifically, we ask: “How likely do you think members of each social group are to get severe COVID-19 or die from COVID-19?” - capturing an indirect social influence rather than a direct influence. It is reasonable to assume that if one believes COVID-19 is dangerous to their family and close friends, they may be more likely to get vaccinated to protect them. In Table 4, we again show an ordinal logistic regression model, where the dependent variable is if the participant is vaccinated or not and a mix of demographic controls and perceived danger are independent variables. In Figure 4, we show the distributions of perceived danger to varying social groups.
Across all social groups, there were no significant differences in perceived danger between vaccinated and unvaccinated participants. While when examining the distributions independently (not controlling for other factors), we see some shifts of perceived danger between vaccinated and unvaccinated groups, particularly for family (visually can be seen in Figure 3a), these differences do not robustly hold. On average, vaccinated participants perceived the danger of COVID-19 between ‘Somewhat unlikely’ to ‘Neither likely nor unlikely’ for all social groups (means ranging from $1.28$ to $2.90$). Similarly, unvaccinated participants perceived the danger of COVID-19 between ‘Somewhat unlikely’ to ‘Neither likely nor unlikely’ for all social groups except close friends (means ranging from $1.97$ to $2.42$).
@@
While to the best of our knowledge, perceived danger of COVID-19 to specific social groups has not been studied, our results are somewhat unexpected when compared with more general findings in the literature. Many factors have been shown to predict COVID-19 risk perceptions, such as personal experience with the virus and hearing about the virus from friends and family dryhurst2020risk . Hearing about the virus through friends and family was shown to be the strongest predictor of perceived risk, where risk includes both risk of self and risk of others contracting the virus dryhurst2020risk . Other non-social factors have also been shown to predict personal risk to COVID-19, including psychological factors and institutional trust dryhurst2020risk ; yildirim2022factor . However, many work related to risk perceptions, particularly studies later or post the COVID-19 pandemic, have focused on risk perceptions of the vaccine rather than the virus lazer2022covid . Perhaps more importantly, it is likely that perceived danger of COVID-19 changed over time, decreasing throughout the pandemic. Our survey reflects views in mid-2021, over a year after the pandemic had started.
3.5 Additional Context: Social Circle Size and Trust
@@
We can further contextualize our findings by approximating the size of and trust in participants social groups. While theoretically, our survey questions measure differences across strong and weak social ties, the strength of the ties between participants and these groups likely changes across individuals. Here we utilize the answers to the questions asked about voluntary conversations participants had within the past week and the county in which participants lived during the pandemic, as described in the methods section.
@@
To better measure differences between the two groups, we perform a Mann-Whitney test on each pair of distributions, finding significant differences between approximate social circle size ($p<0.001$) and the number of individuals trusted ($p<0.001$), but no significant difference was found between the proportion of social circle members trusted ($p=0.9807$). Hence, vaccinated participants interacted with slightly more people than unvaccinated, but trust in those people is no different between the groups. This approximations of the number of people participants interacted with prior to the survey suggest that the vaccinated population interacted with more people than the unvaccinated population did. Theoretically, it makes sense that interacting with less people could impact one’s perceptions of social agreement with COVID-19 preventive measures and COVID-19 itself. In general, this idea is supported in the literature. For example, when studying vaccine messaging strategies, it was found that “messages evoking harm reduction and people you know were more effective in counties where the virus is spreading more quickly. green2021covid
Saliently, there are likely other confounding factors related to location that may have a significant impact on one’s pandemic-related decisions. For example, prior work has demonstrated that local news coverage during 2020 and 2021 was best explained by national trends rather than by local conditions, and the themes of that coverage varied across counties in the U.S. joseph2022local . Further, there is evidence that there were urban-rural differences in COVID-19 behaviors (despite those differences not showing robustly in this study), and that those behaviors changed depending on the news produced locally. Specifically, Kim et al. (2020) showed that “rural residents [were] more likely to engage in social distancing behavior than otherwise similar rural residents if their local news [was] produced in a city that is more impacted by COVID-19” kim2020effect . Other work has shown wide variations in vaccine support across areas in the United States perlis2020covid .
3.6 Additional Context: Open-ended Reasoning